# KAG: Boosting LLMs in Professional Domains via Knowledge Augmented Generation

### v1

@@ 2.5.4 Onepass Inference
As shown in experiment results in Table 8, we draw the following conclusions: (1) OneGen demonstrates efficacy in ${R\rightarrow G}$ task, and joint training of retrieval and generation yields performance gains on the RAG task. The Self-RAG endows LLMs with self-assessment and adaptive retrieval, while OneGen adds self-retrieval. Our method outperforms the original Self-RAG across all datasets, especially achieving improvements of 3.1pt on Pub dataset and 2.8pt on ARC dataset, validating the benefits of joint training. (2) OneGen is highly efficient in training, with instruction-finetuned LLMs showing strong retrieval capabilities with minimal additional tuning. It requires less and lower-quality retrieval data, achieving comparable performance with just 60K noisy samples and incomplete documents, without synthetic data. For more details on the specific implementation process, please refer to
@@
For more details on the specific implementation process, please refer to
@@ 3.1 Experimental Settings
Datasets. To evaluate the effectiveness of the KAG for knowledge-intensive question-answerring task, we perform experiments on 3 widely-used multi-hop QA datasets, including HotpotQA [11], 2WikiMultiHopQA [9], and MuSiQue [10]. For a fair comparison, we follow IRCoT [21] and HippoRAG [6] utilizing 1,000 questions from each validation set and using the retrieval corpus related to selected questions.
@@ 3.2.1 Overall Results
The end-to-end question answering performance is shown in Table 9. Within the RAG framework leveraging ChatGPT-3.5, HippoRAG demonstrates superior performance compared to NativeRAG. HippoRAG employs a human long-term memory strategy that facilitates the continuous integration of knowledge from external documents into LLMs, thereby significantly enhancing question answering capabilities. Additionally, it can be observed that implementing multi-step retrieval in the RAG framework using IRCoT is superior to single-step retrieval. The main reason is that IRCoT can conduct the next search based on the results of the current retrieval, rather than returning all results at once. This also increases the relevance of the retrieved content. Given the substantial operational costs associated with utilizing ChatGPT-3.5, we opted to use the DeepSeek-V2 API as a viable alternative. On average, the performance of the IRCoT + HippoRAG configuration utilizing the DeepSeek-V2 API slightly surpasses that of ChatGPT-3.5. Our constructed framework KAG shows significant performance improvement compared to IRCoT + HippoRAG, with EM increases of 11.5%, 19.8%, and 10.5% on HotpotQA, 2WikiMultiHopQA, and MuSiQue respectively, and F1 improvements of 12.5%, 19.6%, and 12.2%. These advancements in end-to-end performance can largely be attributed to the development of more effective indexing and retrieval libraries within our framework. We evaluate the effectiveness of the single-step retriever and multi-step retriever, with the retrieval performance shown in Table 10. From the experimental results, it is evident that the multi-step retriever generally outperforms the single-step retriever. Analysis reveals that the content retrieved by the single-step retriever exhibits very high similarity, resulting in an inability to use the single-step retrieval outcomes to derive answers for certain data that require reasoning. The multi-step retriever alleviates this issue. Our proposed KAG framework directly utilizes the multi-step retriever and significantly enhances retrieval performance through strategies such as mutual indexing, logical form solving, and knowledge alignment. Compared to IRCoT + HippoRAG, KAG achieves an average Recall@2 improvement of 10.2% and an average Recall@5 improvement of 12.0% on HotpotQA, 2WikiMultiHopQA, and MuSiQue.
@@
KAG (ours) DeepSeek-V2 62.5 76.2 67.8 76.7 36.7 48.7
Table 9: The end-to-end generation performance of different RAG models on three multi-hop question answering datasets. Bold text indicates that the same base model performs best. NativeRAG and HippoRAG use single-step retrieval, while other models employ multi-step retrieval.
Retriever HotpotQA 2WikiMultiHopQA MuSiQue
@@
Table 10: The performance of different retrieval models on three multi-hop question-answering datasets.
3.3 Model Ablation Studies
details
@@ 5.1 DIKW Pyramid
Following the DIKW pyramid theories[29, 30, 31, 32], after data is processed and contextualised, it becomes information, and by integrating information with experience, understanding, and expertise, we gain knowledge. We usually use information extraction technology to obtain information from the original text[33, 34, 35], and obtain knowledge from the information through linking, fusion, analysis, and learning technology[31, 36, 34]. Information and knowledge are a single entity having different forms. There are no unified language to represent data, information and knowledge, RDF/OWL[37] only provides binary representation in the form of triples, and LPG[12] lacks support for knowledge semantics and classification. SPG [^4] Official site of SPG: https://spg.openkg.cn/en-US[38] supports knowledge hierarchy and classification representation, but lacks text context support that is friendly to large language models. Our proposed LLMFriSPG supports hierarchical representation from data to information to knowledge, and also provides reverse context-enhanced mutual indexing.
@@ 5.3 Information Retrieval-based GraphRAG
5.4 KG-based Query and Answering
@@ 6 Conclusion and Future Work
In order to build professional knowledge services in vertical domains, fully activate the capabilities and advantages of symbolic KGs and parameterized LLMs, and at the same time significantly reduce the construction cost of domain KGs, we proposed the KAG framework and try to accelerated its application in professional domains. In this article, we introduce in detail the knowledge accuracy, information completeness and logical rigorous are the key characteristics that professional knowledge services must have. At the same time, we also introduce innovations such as LLMs friendly knowledge representation, mutual index of knowledge structure and text chunks, knowledge alignment by semantic reasoning, logic-form-guided hybrid reasoning&solving and KAG model. Compared with the current most competitive SOTA method, KAG has achieved significant improvements on public data sets such as HotpotQA, 2wiki, musique. We have also conducted case verifications in E-goverment Q&A and E-Health Q&A scenarios of Alipay, further proving the adaptability of the KAG framework in professional domains.
In the future, there is still more work to be explored to continuously reduce the cost of KG construction and improve the interpretability and transparency of reasoning, such as multiple knowledge extraction, knowledge alignment based on OneGraph, domain knowledge injection, large-scale instruction synthesis, illusion suppression of knowledge logic constraints, etc. We will also work in depth with the community organization OpenKG to continue to tackle key technical issues in the collaboration between LLMs and KGs.
@@ 7 Acknowledgements
This work was completed by the AntGroup Knowledge Graph Team, thank you all for your continuous innovation attempts and hard work. This work also received strong support from Professor Huajun Chen, Researcher Wen Zhang of Zhejiang University, and Professor Wenguang Chen of AntGroup Technology Research Institute, thank you all.

### v2

@@ 2.5.4 Onepass Inference
As shown in experiment results in Table 7, we draw the following conclusions: (1) OneGen demonstrates efficacy in ${R\rightarrow G}$ task, and joint training of retrieval and generation yields performance gains on the RAG task. The Self-RAG endows LLMs with self-assessment and adaptive retrieval, while OneGen adds self-retrieval. Our method outperforms the original Self-RAG across all datasets, especially achieving improvements of 3.1pt on Pub dataset and 2.8pt on ARC dataset, validating the benefits of joint training. (2) OneGen is highly efficient in training, with instruction-finetuned LLMs showing strong retrieval capabilities with minimal additional tuning. It requires less and lower-quality retrieval data, achieving comparable performance with just 60K noisy samples and incomplete documents, without synthetic data. For more details on the specific implementation process, please refer to paper[25]
@@ 3.1 Experimental Settings
Datasets. To evaluate the effectiveness of the KAG for knowledge-intensive question-answering task, we perform experiments on 3 widely-used multi-hop QA datasets, including HotpotQA [13], 2WikiMultiHopQA [11], and MuSiQue [12]. For a fair comparison, we follow IRCoT [26] and HippoRAG [10] utilizing 1,000 questions from each validation set and using the retrieval corpus related to selected questions.
@@ 3.2.1 Overall Results
The end-to-end Q&A performance is shown in Table 8. Within the RAG frameworks leveraging ChatGPT-3.5 as backbone model, HippoRAG demonstrates superior performance compared to NativeRAG. HippoRAG employs a human long-term memory strategy that facilitates the continuous integration of knowledge from external documents into LLMs, thereby significantly enhancing Q&A capabilities. However, given the substantial economic costs associated with utilizing ChatGPT-3.5, we opted to use the DeepSeek-V2 API as a viable alternative. On average, the performance of the IRCoT + HippoRAG configuration utilizing the DeepSeek-V2 API slightly surpasses that of ChatGPT-3.5. Our constructed framework KAG shows significant performance improvement compared to IRCoT + HippoRAG, with EM increases of 11.5%, 19.8%, and 10.5% on HotpotQA, 2WikiMultiHopQA, and MuSiQue respectively, and F1 improvements of 12.5%, 19.1%, and 12.2%. These advancements in end-to-end performance can largely be attributed to the development of more effective indexing, knowledge alignment and hybrid solving libraries within our framework. We evaluate the effectiveness of the single-step retriever and multi-step retriever, with the retrieval performance shown in Table 9. From the experimental results, it is evident that the multi-step retriever generally outperforms the single-step retriever. Analysis reveals that the content retrieved by the single-step retriever exhibits very high similarity, resulting in an inability to use the single-step retrieval outcomes to derive answers for certain data that require reasoning. The multi-step retriever alleviates this issue. Our proposed KAG framework directly utilizes the multi-step retriever and significantly enhances retrieval performance through strategies such as mutual-indexing, logical form solving, and knowledge alignment.
@@
KAG w/ $LFS_{ref_{3}}$ DeepSeek-V2 59.8 74.0 66.3 76.1 35.4 48.2
KAG w/ $LFSH_{ref_{3}}$ DeepSeek-V2 62.5 76.2 67.8 76.2 36.7 48.7
Table 8: The end-to-end generation performance of different RAG models on three multi-hop Q&A datasets. The values in bold and underline are the best and second best indicators respectively.
Retriever HotpotQA 2Wiki MuSiQue
@@
Table 9: The performance of different retrieval models on three multi-hop Q&A datasets
3.3 Ablation Studies
The objective of this experiment is to deeply investigate the impact of the knowledge alignment and logic form solver on the final results. We conduct ablation studies for each module by substituting different methods and analyzing the changes in outcomes.
3.3.1 Knowledge Graph Indexing Ablation
In the graph indexing phase, we propose the following two substitution methods:
1) Mutual Indexing Method. As a baseline method of KAG, according to the introduction in Sections 2.1 and 2.2, we use information extraction methods (such as OpenIE) to extract phrases and triples in document chunks, and form the mutual-indexing between graph structure and text chunks according to the hierarchical representation of LLMFriSPG, and then write them into KG storage. We denote this method as M_Indexing.
2) Knowledge Alignment Enhancement. This method uses knowledge alignment to enhance the KG mutual-indexing and the logical form-guided reasoning & retrieval. According to the introduction in Section 2.4, it mainly completes tasks such as the classification of instances and concepts, the prediction of hypernyms/hyponyms of concepts, the completion of the semantic relationships between concepts, the disambiguation and fusion of entities, etc., which enhances the semantic distinction of knowledge and the connectivity between instances, laying a solid foundation for subsequent reasoning and retrieval guided by logical forms. We denote this method as K_Alignment.
3.3.2 Reasoning and Retrieval Ablation
Multi-round Reflection. We adopted the multi-round reflection mechanism from ReSP[19] to assess whether the Logical Form Solver has fully answered the question. If not, supplementary questions are generated for iterative solving until the information in global memory is sufficient. We analyzed the impact of the maximum iteration count $n$ on the results, denoted as $ref_{n}$. If $n=1$, it means that the reflection mechanism is not enabled. In the reasoning and retrieval phase, we design the following three substitution methods:
1) Chunks Retriever. We define KAG’s baseline retrieval strategy with reference to HippoRAG’s[10] retrieval capabilities, with the goal of recalling the top_k chunks that support answering the current question. The Chunk score is calculated by weighting the vector similarity and the personalized pagerank score. We denote this method as $ChunkRetri$, we denote ChunkRetri with n-round reflections as $CR_{ref_{n}}$.
2) Logical Form Solver (Enable Graph Retrieval). Next, we employ a Logical Form Solver for reasoning. This method uses pre-defined logical forms to parse and answer questions. First, it explores the reasoning ability of the KG structure in $KG_{cs}$ and $KG_{fr}$ spaces, focusing on accuracy and rigor in reasoning. Then, it uses supporting_chunks in $RC$ to supplement retrieval when the previous step of reasoning has no results. We denote this method as $LFS_{ref_{n}}$. The parameter $n$ is maximum number of iteration parameter.
3) Logical Form Solver (Enable Hybrid Retrieval). In order to make full use of the mutual-indexing structure between $KG_{fr}$ and $RC$ to further explore the role of KG structure in enhancing chunk retrieval, we modify the $LFS_{ref_{n}}$ by disabling the Graph Retrieval functionality for direct reasoning. Instead, all answers are generated using the Hybrid Retrieval method. This approach enables us to evaluate the contribution of graph retrieval to the performance of reasoning. We denote this method as $LFSH_{ref_{n}}$.
Through the design of this ablation study, we aim to comprehensively and deeply understand the impact of different graph indexing and reasoning methods on the final outcomes, providing strong support for subsequent optimization and improvement.
3.3.3 Experimental Results and Discussion
Graph Index Reasoning HotpotQA 2Wiki MuSiQue
EM F1 EM F1 EM F1
M_Indexing $CR_{ref_{3}}$ 52.4 65.4 48.2 56.0 24.6 36.6
K_Alignment $CR_{ref_{3}}$ 54.7 69.5 62.7 72.5 29.6 41.1
$LFS_{ref_{1}}$ 59.1 73.4 65.2 74.4 31.3 43.4
$LFS_{ref_{3}}$ 59.8 74.0 66.3 76.1 35.4 48.2
$LFSH_{ref_{1}}$ 61.5 76.0 66.0 75.0 33.5 44.3
$LFSH_{ref_{3}}$ 62.5 76.2 67.8 76.2 36.7 48.7
Table 10: The end-to-end generation performance of different model methods on three multi-hop Q&A datasets. The backbone model is DeepSeek-V2 API. As is described in Algorithm 17, $ref_{3}$ represents a maximum of 3 rounds of reflection, and $ref_{1}$ represents a maximum of 1 round, which means that no reflection is introduced.
Graph Index Reasoning HotpotQA 2Wiki MuSiQue
R@2 R@5 R@2 R@5 R@2 R@5
M_Indexing $CR_{ref_{3}}$ 61.5 73.8 54.6 59.7 39.3 52.8
K_Alignment $CR_{ref_{3}}$ 56.3 83.0 66.3 88.1 40.0 62.3
$LFS_{ref_{1}}$ / / / / / /
$LFS_{ref_{3}}$ / / / / / /
$LFSH_{ref_{1}}$ 55.1 85.0 65.9 92.4 36.1 58.4
$LFSH_{ref_{3}}$ 72.7 88.8 65.4 91.9 48.4 65.6
Table 11: The recall performance of different methods across three datasets is presented. The answers to some sub-questions in the $LFS_{ref_{n}}$ method use KG reasoning without recalling supporting chunks, which is not comparable to other methods in terms of recall rate. BackBone model is DeepSeek-V2 API.
Figure 8: Each of the three test datasets comprises 1000 test problems, with 20 tasks processed concurrently and maximum number of iterations $n$ is 3. $CR_{ref_{3}}$ method exhibits the fastest execution, whereas $LFSH_{ref_{3}}$ method is the slowest. Specifically, $CR_{ref_{3}}$ method outperforms $LFSH_{ref_{3}}$ method by 149%, 101%, and 134% across the three datasets. In comparison, on the same dataset, the $LFS_{ref_{3}}$ method outperforms the $LFSH_{ref_{3}}$ method by 13%, 22%, and 18%, respectively, with F1 relative losses of 2.6%, 0.1%, and 1.0%, respectively.
Figure 9: The connectivity of the graph exhibits a notable rightward shift after applying K_Alignment,the distribution changes of 1-hop, 2-hop, and 3-hop neighbors are shown.
The analysis of the experimental outcomes can be approached from the following two perspectives:
1) Knowledge Graph Indexing. As is shown in Table 11, after incorporation Knowledge Alignment into the KG mutual-indexing, the top-5 recall rates of $CR_{ref_{3}}$ improved by 9.2%, 28.4%, and 9.5% respectively, with an average improvement of 15.7%. As shown in Figure 9, after enhancing knowledge alignment, the relation density is significantly increased, and the frequency-outdegree graph is shifted to the right as a whole
  • The 1-hop graph exhibits a notable rightward shift, indicating that the addition of semantic structuring has increased the number of neighbors for each node, thereby enhancing the graph’s density.
  • The 2-hop and 3-hop graphs display an uneven distribution, with sparse regions on the left and denser regions on the right. When comparing before and after K_Alignment, it is evident that the vertices in each dataset have shifted rightward, with the left side becoming more sparse. This suggests that nodes with fewer multi-hop neighbors have gained new neighbors, leading to this observed pattern.
This signifies that the newly added semantic relations effectively enhance graph connectivity, thereby improving document recall rates.
2) Graph Inference Analysis. In terms of recall, $LFSH_{ref_{3}}$ achieves improvements over $CR_{ref_{3}}$ under the same graph index, with increases in top-5 recall rates by 15%, 32.2%, and 12.7%, averaging an improvement of 19.9%. This enhancement can be attributed to two main factors:
  • $LFSH_{ref_{3}}$ decomposes queries into multiple executable steps, with each sub-query retrieving chunks individually. As shown in the time analysis in Figure 8, both $LFSH_{ref_{3}}$ and $LFS_{ref_{3}}$ consume more than twice the time of $LFSH_{ref_{3}}$, indicating that increased computational time is a trade-off for improved recall rates.
  • $LFSH_{ref_{3}}$ not only retrieves chunks but also integrates SPO triples from execution into chunk computation. Compared to $LFSH_{ref_{3}}$, it retrieves additional query-related relationships.
Due to the subgraph-based query answering in $LFS_{ref_{3}}$, it cannot be compared directly in recall rate analysis but can be examined using the F1 metric. In comparison to $LFSH_{ref_{3}}$, $LFS_{ref_{3}}$ answered questions based on the retrieved subgraphs with proportions of 33%, 34%, and 18%,respectively. $LFS_{ref_{3}}$ shows a decrease in the F1 metric by 2.2%, 0.1%, and 0.5%, while the computation time reduces by 12%, 22%, and 18%.
The analysis of the cases with decreased performance reveals that errors or incomplete SPOs during the construction phase lead to incorrect sub-query answers, resulting in wrong final answers. This will be detailed in the case study. The reduction in computation time is primarily due to the more efficient retrieval of SPOs compared to document chunks.
In industrial applications, computation time is a crucial metric. Although $LFS_{ref_{n}}$ may introduce some errors, these can be improved through graph correction and completion. It is noteworthy that in the current experiments, the slight decrease in metrics has been traded off for reduced computation time, which we consider a feasible direction for industrial implementation.
For analyze the impact of the maximum number of iterations parameter $n$ on the results, $LFS_{ref_{1}}$ compared to $LFS_{ref_{3}}$ , the F1 scores decreased by 0.6%, 1.6%, and 4.8%, respectively. Based on the experiments of $LFS_{ref_{3}}$, the proportions for an iteration count of 1 were analyzed to be 97.2%, 94.8%, and 87.9%; $LFSH_{ref_{1}}$ compared to $LFSH_{ref_{3}}$ , the F1 scores decreased by 0.2%, 1.2%, and 4.4%, respectively. Based on the experiments of $LFSH_{ref_{3}}$, the proportions for an iteration count of 1 were analyzed to be 98.3%, 95.2%, and 84.1%; showing a positive correlation with the F1 score reduction. Table 13 provides a detailed analysis of the effect of iteration rounds on the solution of the final answer. Increasing the maximum number of iterations parameter facilitates the re-planning of existing information when $LFS_{ref_{n}}$ is unable to complete the solution, thereby addressing some unsolvable case.
@@ 5.1 DIKW Pyramid
Following the DIKW pyramid theories[34, 35, 36, 37], after data is processed and contextualised, it becomes information, and by integrating information with experience, understanding, and expertise, we gain knowledge. We usually use information extraction technology to obtain information from the original text[38, 39, 40], and obtain knowledge from the information through linking, fusion, analysis, and learning technology[36, 41, 39]. Information and knowledge are a single entity having different forms. There are no unified language to represent data, information and knowledge, RDF/OWL[42] only provides binary representation in the form of triples, and LPG[14] lacks support for knowledge semantics and classification. SPG [^4] Official site of SPG: https://spg.openkg.cn/en-US[43] supports knowledge hierarchy and classification representation, but lacks text context support that is friendly to large language models. Our proposed LLMFriSPG supports hierarchical representation from data to information to knowledge, and also provides reverse context-enhanced mutual-indexing.
@@ 5.3 Information Retrieval-based GraphRAG
5.4 KG-based Question and Answering
@@ 5.5 Bidirectional-enhancement of LLMs and KGs
6 Limitations
In this article, we have proven the adaptability of the KAG framework in Q&A scenarios in vertical and open domains. However, the currently developed version of OpenSPG-KAG 0.5 still has major limitations that need to be continuously overcome, such as:
Implementing our framework requires multiple LLM calls during the construction and solving phases. A substantial number of intermediate tokens required to be generated during the planning stage to facilitate the breakdown of sub-problems and symbolic representation, this leads to computational and economic overhead, as illustrated in Table 14, where the problem decomposition not only outputs sub-problems but also logical functions, resulting in approximately twice as many generated tokens compared to merely decomposing the sub-problems. Meanwhile, currently, all model invocations within the KAG framework, including entity recognition, relation extraction, relation recall, and standardization, rely on large models. This multitude of models significantly increases the overall runtime. In future domain-specific implementations, tasks like relation recall, entity recognition, and standardization could be substituted with smaller, domain-specific models to enhance operational efficiency.
The ability to decompose and plan for complex problems requires a high level of capability. Currently, this is implemented using LLMs, but planning for complex issues remains a significant challenge. For instance, when the task is to compare who is older, the problem should be decomposed into comparing who was born earlier. Directly asking for age is not appropriate, as they are deceased, and "what is the age" refers to the age at death, which doesn’t indicate who is older. Decomposing and planning complex problems necessitates ensuring the model’s accuracy, stability, and solvability in problem decomposition and planning. The current version of the KAG framework does not yet address optimizations in these areas. We will further explore how pre-training, SFT, and COT strategies can improve the model’s adaptability to logical forms and its planning and reasoning capabilities.
Question: Which film has the director who is older, God’S Gift To Women or Aldri Annet Enn Bråk?
Q1: Which director directed the film God’S Gift To Women? A1: Michael Curtiz
Q2: Which director directed the film Aldri Annet Enn Bråk? A2: Edith Carlmar
Q3: What is the age of the director of God’S Gift To Women? A3: 74 years old. Michael Curtiz (December 24, 1886 to April 11, 1962)…
Q4: What is the age of the director of Aldri Annet Enn Bråk? A4: 91 years old. Edith Carlmar (Edith Mary Johanne Mathiesen) (15 November 1911 to 17 May 2003) …
Q5: Compare the ages of the two directors to determine which one is older. A5: Edith Carlmar is older. Actually, Michael Curtiz was born earlier.
OpenIE significantly lowers the threshold for building KGs, but it also obviously increases the technical challenges of knowledge alignment. Although the experiments in this article have shown that the accuracy and connectivity of extracted knowledge can be improved through knowledge alignment. However, there are still more technical challenges waiting to be overcome, such as optimizing the accuracy of multiple-knowledge(such as events, rules, pipeline, etc.) extraction and the consistency of multiple rounds of extraction. In addition, schema-constraint knowledge extraction based on the experience of domain experts is also a key way to obtain rigorous domain knowledge, although the labor cost is high. These two methods should be applied collaboratively to better balance the requirements of vertical scenarios for the rigor of complex decision-making and the convenience of information retrieval. For instance, when extracting team members from multiple texts and asked about the total number of team members, a comprehensive extraction is crucial for providing an accurate answer based on the structured search results. Incorrect extractions also impair response accuracy.
@@ 7 Conclusion and Future Work
In order to build professional knowledge services in vertical domains, fully activate the capabilities and advantages of symbolic KGs and parameterized LLMs, and at the same time significantly reduce the construction cost of domain KGs, we proposed the KAG framework and try to accelerated its application in professional domains. In this article, we introduce in detail the knowledge accuracy, information completeness and logical rigorous are the key characteristics that professional knowledge services must have. At the same time, we also introduce innovations such as LLMs friendly knowledge representation, mutual-indexing of knowledge structure and text chunks, knowledge alignment by semantic reasoning, logic-form-guided hybrid reasoning&solving and KAG model. Compared with the current most competitive SOTA method, KAG has achieved significant improvements on public data sets such as HotpotQA, 2wiki, musique. We have also conducted case verifications in E-goverment Q&A and E-Health Q&A scenarios of Alipay, further proving the adaptability of the KAG framework in professional domains.
In the future, there is still more work to be explored to continuously reduce the cost of KG construction and improve the interpretability and transparency of reasoning, such as multiple knowledge extraction, knowledge alignment based on OneGraph, domain knowledge injection, large-scale instruction synthesis, illusion suppression of knowledge logic constraints, etc.
This study does not encompass the enhancement of models for decomposing and planning complex problems, which remains a significant area for future research. In future work, KAG can be employed as a reward model to provide feedback and assess the model’s accuracy, stability, and solvability through the execution of planning results, thereby enhancing the capabilities of planning models.
We will also work in depth with the community organization OpenKG to continue to tackle key technical issues in the collaboration between LLMs and KGs.
@@ 8 Acknowledgements
This work was completed by the AntGroup Knowledge Graph Team, in addition to the authors in the list, other contributors include Yuxiao He, Deng Zhao, Xiaodong Yan, Dong Han, Fanzhuang Meng, Yang Lv, Zhiying Yin, etc, thank you all for your continuous innovation attempts and hard work. This work also received strong support from Professor Huajun Chen, Researcher Wen Zhang of Zhejiang University, and Professor Wenguang Chen of AntGroup Technology Research Institute, thank you all.
Appendix A Example of KAG Solver
Round One
Initial Question: How many times did the plague occur in the birth place of Concerto in C Major Op 3 6’s composer?
Step1: Who is the composer of Concerto in C Major Op 3 6?.
Logical Form: Retrieval(s=s1:Work[C Major Op 3 6],p=p1:composer,o=o1:Person)
Graph Retrieval: (antonio vivaldi, composed concerto in, c major op 3 6)
Hybird Retrieval: not executed
A1: Antonio Vivaldi
Step2: What is the birthplace of this composer?
Logical Form: Retrieval(s=o1,p=p2:brithPlace,o=o2:Place)
Graph Retrieval: not found
Hybird Retrieval:
doc1:Antonio Vivaldi composed a mandolin concerto (Concerto in C major Op.3 6) …
doc2:Antonio Lucio Vivaldi (Italian: … Born in Venice
A2: Venice
Step3: How many times did the plague occur in this birthplace?
Logical Form: Retrieval(s=o2,p=p3:plagueOccurrence,o=o3:Plague)
Graph Retrieval: not found
Hybird Retrieval:
doc1:Antonio Vivaldi composed a mandolin concerto (Concerto in C major Op.3 6) …
doc2:The Black Death, also known as the Great Plague.. in Europe from 1347 to 1351..
A3: I don’t know
Step4: Count the number of plague occurrences.
Logical Form: Count(o3)->nums
Graph Retrieval: not found
Hybird Retrieval: not found
A4: I don’t know
Round Two
Rewrite Question: What specific records or historical accounts detail the number of plague occurrences in Venice during the mid-14th century?
Step1: What specific records or historical accounts detail the number of plague occurrences in Venice during the mid-14th century?
Logical Form: Retrieval(s=s1:City[Venice],p=p3:plagueOccurrencesInMid14thCentury,o=o1:Times)
Graph Retrieval: not found
Hybird Retrieval:
doc1:In 1466, perhaps 40,000 people died of the plague…Plague occurred in Venice 22 times between 1361 and 1528
doc2:The Black Death, also known as the Great Plague…
A1: The plague occurred in Venice 22 times between 1361 and 1528. The 1576–77 plague killed 50,000, almost a third of the population.
Final Answer
Question: How many times did the plague occur in the birth place of Concerto in C Major Op 3 6’s composer?
Step1: What specific records or historical accounts detail the number of plague occurrences in Venice during the mid-14th century?
A: 22 times
Table 13: An example of using logical-from to guide question planning, reasoning, retrieval, and answer generation, and using multiple rounds of reflection to rephrase questions.
Appendix B Example of Logical form Reasoner
Numerical Reasoning
Logical Form: Retrieval(s=s1:Player[Cristiano Ronaldo], p=p1:playedFor, o=o1:SportsTeam, p.PlayedForInYear=2011)
Step2: Determine the Foundation Years of Each Identified Team.
Logical Form: Retrieval(s=o1, p=p2:foundationYear, o=o2:Year)
Step3: Which team was founded last?
Logical Form: Sort(set=o1, orderby=o2, direction=max, limit=1)
Step1: What is the sum of 30 + 6 ?
Logical Form: math1 = Math(30+6)
Step2: Who is the founder of Tesla?
Logical Form: Retrieval(s=s2:Company[Tesla], p=p2:founder, o=o2)
Step3: In which year was the founder of Tesla born?
Logical Form: Retrieval(s=o2, p=p3:yearOfBirth, o=o3)
Step4: How old will the founder of Tesla be in the year 2027?
Logical Form: math4 = Math(2027-o3)
Step5: What is the sum of math1 and math4?
Logical Form: math5 = Math(math1+math4)
Logical Reasoning
question: Find a picture containing vegetables or fruits.
Step1: Find pictures containing vegetables.
Logical Form: Retrieval(s=s1:Image, p=p2:contains, o=o1:Vegetables)
Step2: Find pictures containing fruits.
Action2: Retrieval(s=s2:Image, p=p2:contains, o=o2:Fruits)
Step3: Output s1, s2.
Logical Form: Output(s1, s2)
question: Find a picture containing vegetables and fruits.
Step1: Find pictures containing vegetables.
Logical Form: Retrieval(s=s1:Image, p=p2:contains, o=o1:Vegetables)
Step2: Find pictures containing fruits.
Logical Form: Retrieval(s=s1, p=p2:contains, o=o2:Fruits)
Step3: Output s1.
Logical Form: Output(s1)
Semantic Deduce
Step1: What documents are required to apply for a passport?
Logical Form: Retrieval(s=s1:Event[apply for a passport], p=p1:support_chunks, o=o1:Chunk)
Step2: Does this set of documents include the original identity card?
Logical Form: Deduce(left=o1, right=the original identity card, op=entailment)
Table 14: The cases of reasoning with logical form